OpenStudy (anonymous):

An object moves along a coordinate line so that its position "s" satisfies the eqn.: s = 4t^2-4t+8. The unit "s" is measured in cm, and "t" is measured in seconds. 1.) Determine velocity of object when t = 1 2.) Determine velocity of object when t = 6 3.) When the velocity is 0, how many seconds have passed?

OpenStudy (stamp):

Is this a calculus class, or a calculus based physics class?

OpenStudy (anonymous):

After the questions we have been getting lately, I was beginning to wonder the same thing! (The class is Calculus I)

OpenStudy (stamp):

Calculus and physics are the same thing pretty much. Newton was a physicist who developed calculus to aid in his physics. Since you are in Calculus, you know that the derivative of position is velocity. \[s(t)=4t^2-4t+8\]\[s'(t)=v(t)\]Find s'(t)

OpenStudy (stamp):

Velocity as the derivative of position makes sense, think how your change in position over change in time\[\Delta position/\Delta time=distance/time=km/h,\ mi/h, etc.\]

OpenStudy (anonymous):

Thank You for your help stamp! I will try to apply your wisdom to the best of my ability.

OpenStudy (stamp):

What did you get for s'(t)=v(t)?

OpenStudy (anonymous):

(Still working on it!)

OpenStudy (stamp):

@mylittlepony are you using the power rule? What methods have you learned in lecture for finding a derivative of a function.

OpenStudy (anonymous):

Using the power rule, I took the derivative of the position function (as you advised) to yield an equation for velocity. Position Function: 4t^2-4t+8 = s Derivative: 8t-4 = s/t 1.) t=1, (s/t)=4 2.) t=6, (s/t)=44 3.) t = 1/2, t/s=0 Thank You so much Stamp!!! (I would have never gotten this question w/o you!)

OpenStudy (stamp):

\[s(t)=4t^2-4t+8\]\[s'(t)=v(t)=8t-4=4(2t-1)\]\[(a).\ Find\ v(1).\]\[v(1)=8(1)-4=4\ cm/s\]\[(b).\ Find\ v(6)\]\[v(6)=8(6)-1=45\ cm/s\]\[(c).\ Find\ t\ when\ v(t)=0\]\[0=8t-4\]\[4(2t-1)=0\]\[(2t-1)=0\]\[v(t)=0\ when\ t=1/2\ s\]

OpenStudy (stamp):

I made an error on (b), it should be\[v(t)=8t-4\]\[v(6)=8(6)-4=44\ cm/s\]